A trader brought a television set for $47500.00. He allowed a customer a discount of 5% off marked price. If the trader still made a profit of 25%, calculate.
(a) the traders selling price to the customer
(b) the marked price

Found 2 solutions by Theo, MathTherapy:
Answer by Theo(13342)

cost to trader = 47500.00
discount to customer = 5%.
profit = 25% based on discounted selling price.
selling price minus cost, divided by cost = profit.
s = selling price.
(s - 47500)/47500 = .25
multiply both sides of this equation by 47500 to get:
s - 47500 = .25 * 47500
add 47500 to both sides of the equation to get:
s = .25 * 47500 + 47500 = 59375.
the cost was 47500 and the selling price was 59375.
(59375 - 47500) / 47500 = .25
the selling price was a 5% discount over the marked price.
m = markup price
s = m - .05 * m = .95 * m
solve for m to get:
m = s / .95 = 59375 / .95 = 62500.
the markup price was 62500.
the discount was 5% * 62500 = 3125
the selling price was 62500 minus 3125 = 59375.
the cost was 47500
the profit was (59375 - 47500) / 47500 = .25 = 25%
answers to your questions are:
If the trader still made a profit of 25%, calculate.
(a) the traders selling price to the customer = 59375
(b) the marked price = 62500

Answer by MathTherapy(10552)

A trader brought a television set for $47500.00. He allowed a customer a discount of 5% off marked price. If the trader still made a profit of 25%, calculate.
(a) the traders selling price to the customer
(b) the marked price
Cost: $47,500
Let marked price be M
Then selling price = .95M, since itw's sold at 5% off marked price
As discounted/selling price resulted in a 25% profit we then have: .95M = 1.25(47,500)

                                             Marked price, or 

                                             Selling price: .95($62,500) = $59,375